### v3

# Enabling AI Scientists to Recognize Innovation: A Domain-Agnostic Algorithm for Assessing Novelty

Abstract
As large language models (LLMs) demonstrate increasing capability in generating research ideas, automating novelty evaluation has become a critical challenge for AI-driven scientific discovery. This paper presents Relative Neighbor Density (RND), a domain-agnostic algorithm for novelty assessment in research ideas that overcomes the limitations of existing approaches by comparing an idea’s local density with its adjacent neighbors’ densities. We first developed a scalable methodology to create test set without expert labeling, addressing a fundamental challenge in novelty assessment. Using these test sets, we demonstrate that our RND algorithm achieves state-of-the-art (SOTA) performance in computer science (AUROC=0.820) and biomedical research (AUROC=0.765) domains. Most significantly, while SOTA models like Sonnet-3.7 and existing metrics show domain-specific performance degradation, RND maintains consistent accuracies across domains by its domain-invariant property, outperforming all benchmarks by a substantial margin (0.795 v.s. 0.597) on cross-domain evaluation. These results validate RND as a generalizable solution for automated novelty assessment in scientific research.
footnotetext: Mingxuan CUI did this work during his internship at Yidu Technology.
1 Introduction
As AI research advances toward AGI, automating scientific discovery processes becomes increasingly important. While large language models (LLMs) now demonstrate capability in generating research ideas, such as AI scientist (Lu et al. (2024)), a critical challenge remains in how to pick pearls from those generated ideas, i.e., how to reliably evaluate these ideas for novelty. Traditionally, novelty in research ideas has been assessed through expert evaluation, where domain specialists judge originality based on their knowledge and experience. However, such assessments are inherently subjective, time-consuming, and inconsistent across evaluators. Automated methods for novelty assessment are therefore crucial for identifying truly innovative directions, particularly as AI systems begin to participate more actively in the research process.
Existing approaches primarily fall into two categories: (1) leveraging large language models (LLMs) as judges and (2) using absolute local density-based novelty metrics. LLM-based approaches include Swiss system tournaments (Si et al. (2024); Hu et al. (2024)) and augmented judgment using NeurIPS review guidelines and Semantic Scholar API (Lu et al. (2024); Su et al. (2024)). Alternatively, absolute local density metrics measure novelty through Euclidean distances between embeddings, as demonstrated by Su et al.’s Overall Novelty (ON) metric (Su et al. (2024)), which combines Historical Dissimilarity, Contemporary Dissimilarity, and citation-based Contemporary Impact, showing strong correlation with human-labeled novelty.
Despite these advances, significant challenges remain. First of all, in studies of any existing approach, the validation methodologies typically rely on small, domain-specific test sets with human-labeled data that quickly become outdated as research advances, limiting scalability and long-term accuracy. Additionally, the LLM-based judgment exhibits sensitivity to input perturbations (Zhuo et al. (2023); Singh et al. (2024)), potentially yielding inconsistent novelty ratings for semantically identical ideas. Finally, absolute local density metrics suffer from arbitrary parameter choices and lack generalizability across research domains with varying citation patterns and publication velocities.
To address these challenges, we establish comprehensive semantic embedding databases for novelty assessment, incorporating publications from two distinct domains: Pubmed, the leading biomedical literature search engine with nearly 36 million articles (Jin et al. (2024)), and Arxiv, which contains more than 2.3 million scholarly articles across eight subject areas (Cornell Tech (2023)). Based on these resources, we develop a methodology to create test sets with trustworthy novelty labels without requiring expert manual labeling.
Furthermore, we propose the Relative Neighbor Density (RND) algorithm, which measures novelty by analyzing the distribution patterns of semantic neighbors rather than simple absolute local density (Figure 1). This approach proves more reliable than LLM-based judgments and more generalizable than existing absolute local density-based metrics across different research domains. Our extensive evaluations using test sets from computer science, biomedical science, and cross-domain contexts demonstrate that our proposed algorithm maintains accuracy within specific domains while scaling effectively across diverse research areas.
Our main contributions are:
  • A novel neighbor density-based Relative Neighbor Density (RND) algorithm for assessing research idea novelty that is robust across domains, which holds domain-invariant property
  • A scalable methodology for validating novelty metrics without expert labeling
  • Comprehensive evaluations comparing SOTA reasoning models, LLMs and algorithms for assessing novelty across multiple research domains
Figure 1: Illustration of the Relative Neighbor Density (RND) algorithm. A1/B1: In this step, both the given idea (triangle in A1 and pentagon in B1) and all existing literature in a given research domain are represented in a semantic embedding space. The $P$ nearest neighbors (A/B/C or A’/B’/C’) of the given idea are identified. Then, for each of these neighbors, the neighbor density is computed by identifying $Q$ nearest surrounding neighbors (neighbor sets s1-s3 in A1 and s1’-s3’ in B1). A2/B2: The neighbor densities of the $P$ closest pieces of literature and the given idea are sorted. The RND score of the given idea is determined based on its relative rank among these neighbor densities.
2 Related Works
2.1 LLMs for Novelty Assessment
Recent work has demonstrated promising results in using LLMs as autonomous judges for research novelty. Si et al. (Si et al. (2024)) evaluated this approach using ICLR submissions, converting them into standardized project proposals and conducting pairwise comparisons between accepted and rejected papers(Table LABEL:tab:Standardized_Project_Proposals_prompt). Their Swiss tournament system iteratively paired proposals based on accumulated scores, with Claude-3.5-Sonnet achieving 71.4% accuracy in predicting paper acceptance. As a control measure, they included human expert reranking, which revealed notable discrepancies between automated and human judgments.
Lu et al. (Lu et al. (2024)) expanded this concept with their AI Scientist framework, integrating idea generation, evaluation, and refinement. Their system employs chain-of-thought prompting and external knowledge retrieval via Semantic Scholar API to enhance assessment quality. While showing promise in matching human-level performance, these LLM-based approaches face fundamental challenges in reliability and consistency, as highlighted by studies showing their sensitivity to input variations (Zhuo et al. (2023); Singh et al. (2024)).
2.2 Absolute Local Density-based Metrics
An alternative approach focuses on semantic local density to evaluate novelty. Su et al. (Su et al. (2024)) used the Historical Dissimilarity (HD), which is the average Euclidean distance between the generated abstract embedding and embeddings of the 5 most similar abstracts in the historical literature base. We denote it as "Absolute Local Density" because the average distance is a metric of local density, and they use the value of density directly.
In addition to HD, they also developed Overall Novelty (ON) as below, where CI is Contemporary Impact, the average citation count of the top 5 most similar abstracts in contemporary literature base; CD is Contemporary Dissimilarity, calculated by same algorithm in contemporary literature base.
$ON=\frac{HD\times CI}{CD}$ (1)
The main challenge of local density-based metric is that the density values vary across different domains. In case of evaluating ideas from mixed research domains, the variance would cause a severe degrade in the accuracy.
2.3 Validating Novelty Metrics
A critical limitation in existing research is the lack of scalable validation methodologies. Current approaches (Lu et al. (2024); Su et al. (2024); Hu et al. (2024); Si et al. (2024)) typically rely on small-size literature database and manually labeled test set created by domain experts. For instance, Su et al. (Su et al. (2024)) constructed an "ecosystem" for computer science (CS) using information extracted from 85,217 papers—a dataset that represents only a small fraction of the CS literature available on platforms like arXiv. While their analysis demonstrated promising correlations between novelty scores and human labels across 100 manually evaluated abstracts, the methodology’s reliance on domain-specific expertise significantly constrains its generalizability, which is echoed by reviewer’s comments that only validating CS is "relatively simple" (Openreview Reviewer 2X9t (2025)).
3 Method
3.1 Problem Description
Given a set of ideas $I$,
$I=\{{idea}_{i}\},i\in[1,N]$ (2)
Where ${idea}_{i}$ is a sequence of words or characters in nature language. $N\geq 1$ represents the number of ideas whose novelty needs to be assessed.
The objective is to design a mapping $\mathscr{F}$ from idea space to a score in real value space
$F({idea}_{i})={score}_{i},\quad\text{where}\quad{idea}_{i}\in I,\,{score}_{i}\in\mathbb{R}$ (3)
The novelty score score should be monotonic, meaning that for any two ideas $\text{idea}_{i}$ and $\text{idea}_{j}$, if $\text{idea}_{i}$ is more novel than $\text{idea}_{j}$, then their corresponding scores must satisfy:
$\forall\,\text{idea}_{i},\text{idea}_{j}\in I,\quad\text{idea}_{i}\succ\text{idea}_{j}\Rightarrow F(\text{idea}_{i})>F(\text{idea}_{j})$ (4)
where $\text{idea}_{i}\succ\text{idea}_{j}$ denotes that $\text{idea}_{i}$ is considered more novel than $\text{idea}_{j}$ based on a given novelty criterion.
3.2 Semantic Embedding & Literature Database
Each published literature’s abstract, which is also a sequence of words or characters in natural language, is denoted as $a_{j}$.
The semantic embedding model is a mapping function $\mathscr{G}$, which maps ideas and abstracts into embedding vectors:
$\mathscr{G}(\text{idea}_{i})=\mathbf{v}_{i},\quad\text{where}\quad\mathbf{v}_{i}\in\mathbb{R}^{\text{dims}},$ (5)
$\mathscr{G}(a_{j})=\mathbf{v}_{j},\quad\text{where}\quad\mathbf{v}_{j}\in\mathbb{R}^{\text{dims}}$ (6)
Thus, the preprocessed literature semantic database is represented as a set $A$:
$A=\{(a_{j},\mathbf{v}_{j})\mid j\in[1,M]\}$ (7)
We collected 36 million academic articles from the PubMed Download API (National Library of Medicine (2025)) and 2.6 million papers from the ArXiv dataset (Cornell University (2025)). Among all fetched documents, only those with both a non-empty title and abstract were considered valid for the experiment, resulting in 25,360,114 papers from PubMed and 2,643,057 papers from ArXiv.
For each paper, two semantic embedding vectors were generated—one from its title and another from its abstract—using the M3-Embedding model Chen et al. (2024). The embedding vector dimension, denoted as dimsdims, is 1024. All texts and embedding vectors were stored in Elasticsearch Version 8 for efficient retrieval.
3.3 Algorithm
For each idea $\text{idea}_{i}$ and its embedding $\mathbf{v}_{i}$, we first find its $P$ nearest neighbors using $k$-Nearest Neighbors (KNN) search:
$\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{P}\}=KNN(\mathbf{v}_{i},P,A)$ (8)
where $\mathbf{v}_{j}$ is the $j$-th nearest neighbor of $\mathbf{v}_{i}$, $j\in[1,P]$.
For the idea itself $\mathbf{v}_{i}$ and each neighbor $\mathbf{v}_{j}$, the neighbor density (ND) is defined as below
$ND=\frac{1}{Q}\sum^{Q}_{k=1}d(\mathbf{v},\mathbf{v}_{k})$ (9)
where $\mathbf{v}_{k}$ denotes the $k$-th nearest article’s embedding vector in the literature corpus $A$ and $d(\cdot,\cdot)$ is the cosine distance between two vectors.
We define the set $S_{i}$ that contains the neighbor density values of $\text{idea}_{i}$’s neighbors:
$S_{i}=\{ND_{j}\mid j\in[1,P]\}$ (10)
Finally, we compute the novelty score $\text{score}_{i}$ for $\text{idea}_{i}$ as:
$\text{score}_{i}=\frac{\left|\{ND\in S_{i}\mid ND\leq ND_{i}\}\right|}{|S_{i}|}\times 100$ (11)
In Appendix A, the pseudocode for ND calculation is presented in Algorithm 1; and the pseudocode of complete algorithm is provided in Algorithm 2.
In this algorithm, the selection of values of $P$ and $Q$ is a trade-off between reliability of estimation and other cost. Refer to Appendix A for mathematical analysis and Section 4.3.1 for empirical experimental results. Based on analysis and empirical experiments, we set $P=100$ and $Q=50$.
3.4 Validation without Human Labeling
As a novelty evaluation algorithm, the most challenging point in past research is to find a reliable labeled test set to evaluate the algorithm. Therefore, we propose a new method to construct a convincing test set instead of relying on human experts to annotate it.
For the positive samples (a.k.a novel ideas) in the test set, we select recent articles from top journals or conferences. For the negative samples (a.k.a. non-novel ideas), highly cited articles published before the last few years were selected, also from the research domain’s top journals or conferences. The fundamental principles behind such methodology were: high-quality novel ideas are more likely to be published in recent issues and top journals or conferences; while after time passes, the at-the-time novel ideas were more likely to attract attention and related works, thus become non-innovative at present.
4 Result
4.1 Experiment Setup
4.1.1 Test Set
We have two test sets: NeurIPS, which represents the most advanced research results in the field of computer science, and Nature Medicine, which represents the most cutting-edge papers in the medical field. The sample year distribution of the test sets can be found in Table 1
NeurIPS test set: The initial corpus consists of papers that are Accept (oral) or Accept (spotlight) by Program Chairs at the 2024 NeurIPS conference, which represents the latest research results in computer science. Furthermore, we select articles from the initial corpus that explicitly mention that the papers have obvious novelty in the comments of Program Chairs to form the positive samples of the NeurIPS test set. The comments and decision information of Program Chairs can be obtained on the OpenReview.net website. At the same time, we use the Semantic Scholar API to obtain the 99 most cited papers published in the NeurIPS conference from 2015 to 2020 to form the negative samples of the test set. The titles of all samples are presented in Table LABEL:tab:NIPS_Dataset
Nature Medicine test set: The positive samples of the Nature Medicine test set consist of articles classified as "Article" type, published in Nature Medicine from August 2024 to February 2025, according to the classification on the nature.com website. Articles related to phase 2 or phase 3 trials were excluded. And we used the same method as the negative samples of the NeurIPS test set to obtain 99 articles of Nature Medicine with the highest citation count in the past 15 years as negative samples of the test set. The titles of all samples are presented in Table LABEL:tab:Nature_Medicine_Dataset
Mixed Test set: The mixed test set is the combination of the negative samples from NeurIPS and positive samples from Nature Medicine.
Table 1: Count of Data in Different Time Ranges for NeurIPS and Nature Medicine Test Sets. Positive: novel samples, Negative: non-novel samples.
Test set Label Count
total 2024-2025 2019-2023 2014-2018 -2014
NeurIPS Positive 80 80 0 0 0
Negative 99 0 31 68 0
Nature Medicine Positive 66 66 0 0 0
Negative 99 0 29 32 38
4.1.2 Baseline
To evaluate our algorithm, we selected all existing novelty assessment algorithms as baselines, categorized into two groups: LLM-based and non-LLM-based. Non-LLM-based algorithms, including Relative Neighbor Density(Ours), Historical Dissimilarity(HD), and Overall Novelty(ON), rely solely on literature search and mathematical calculations. Since the output of the literature search for the same query remains consistent, we conducted a single test to assess the algorithm’s performance. In contrast, for LLM-based algorithms, due to the inherent variability of LLM outputs, we ran three tests for each algorithm, calculated the average result, and included the standard deviation in the results.
For all algorithms tested, we use the abstracts of the papers as "ideas" for testing.
Historical Dissimilarity (HD) & Overall Novelty (ON): Both HD and ON are described in Section 2.2. During calculation, the historical database contains papers from 2011 to 2021, and the contemporary database contains papers from 2021 to 2025.
LLM + literature search: Provide LLM with the titles and abstracts of the 10 most relevant papers to the given idea. The model then assesses whether the core concepts of these papers significantly overlap with the idea(Table LABEL:tab:LLM_with_paper_prompt). Sonnet-3.7 (Anthropic (2025)), Deepseek-r1 (Guo et al. (2025)) and GPT-4o (Hurst et al. (2024)) were selected as the LLMs to be tested.
LLM with guideline: Utilize the NeurIPS 2024 review guidelines to assist LLM in evaluating the novelty of ideas(Table LABEL:tab:LLM_with_guideline_prompt). Sonnet-3.7 is selected as the LLM to be tested.
LLM with tournament: First, the idea is transformed into the Standardized Project Proposal format(Table LABEL:tab:Standardized_Project_Proposals_prompt). Next, the novelty of all standardized ideas is assessed using the Swiss tournament method, where ideas are iteratively compared in a structured competition. Finally, each idea is assigned a score based on the number of wins it accumulates throughout the tournament. Sonnet-3.7 is selected as the LLM to be tested.
4.2 Accuracy Evaluation
Table 2: Validation of Different methods, measured by AUROC. HD: Historical Dissimilarity (section 2.2). ON: Overall Novelty (section 2.2). LLM + literature search: supplementing LLM with 10 relevant papers, which were searched by idea’s embedding from our literature database using semantic embedding. LLM with guideline: using NeurIPS 2024 review guideline to help LLM judge the novelty of ideas, which is not applicable to Nature Medicine. Therefore, the results of Nature Medicine and Mixed are marked as not applicable. LLM with tournament: a Swiss system tournament design to evaluate ideas by using LLM as judge.
Model NeurIPS Nature Medicine Mixed
Relative Neighbor Density(Ours) 0.820 0.765 0.795
Absolute Local Density HD 0.856 0.699 0.362
ON 0.584 0.544 0.456
LLM + literature search Sonnet-3.7 $0.813\pm 0.01$ $0.616\pm 0.006$ $0.597\pm 0.004$
Deepseek-r1 $0.710\pm 0.027$ $0.673\pm 0.025$ $0.596\pm 0.049$
GPT-4o $0.567\pm 0.008$ $0.545\pm 0.02$ $0.522\pm 0.022$
Sonnet-3.7 with guideline $0.546\pm 0.035$ NaN NaN
Sonnet-3.7 with tournament $0.496\pm 0.001$ $0.503\pm 0.005$ $0.501\pm 0.004$
As shown in Table 2, our enhanced neighbor density-based novelty measurement algorithm outperforms all baseline models on the Nature Medicine and Mixed test sets, and performs strongly on the NeurIPS test set.
By comparing the results of various LLM-related algorithms, we find that both Sonnet-3.7 with guideline and Sonnet-3.7 with tournament show limited external knowledge input, leading to inaccurate novelty judgments. In contrast, the LLM + literature search method, which incorporates the 10 most relevant papers based on semantic embedding, significantly improves accuracy. This method shows a higher AUROC in computer science than in biomedicine, highlighting the impact of the model’s internal knowledge on judgment outcomes.
Additionally, Sonnet-3.7 and Deepseek-r1 outperform GPT-4o, suggesting that external knowledge enhances the performance of the inference model compared to autoregressive models.
We also observed that the Historical Dissimilarity (HD) metric performed similarly to our method on the Nature Medicine and NeurIPS test sets. However, on the Mixed test set, there was a significant disparity, with our method achieving an AUROC of 0.795, while HD reached only 0.362. This difference reflects the limited generalization ability of HD across domains. In contrast, the distributions of our method’s scores are consistent across all test sets, indicating its robust cross-domain evaluation capability, making it applicable to various research fields.
Figure 2: Comparison of HD & Our score distributions in different domains. 1: In the right panel, the upper and lower bounds of the score exceeded the actual score range ($[0,100]$) because of linear interpolation. 2: to make the horizontal axis comparable, we scaled the Historical Dissimilarity scores by $\times 100$.
4.3 Sensitivity Study
4.3.1 Sensitivity of Hyper-Parameter
As illustrated in the left panel of Figure 3, the AUROC on each test set increases as $P$ grows. However, when $P>50$, the improvement in AUROC becomes marginal compared to the significant gain observed when increasing $P$ from 10 to 50. This suggests that the marginal benefit of further increasing $P$ diminishes while simultaneously incurring substantial computational costs, given that the algorithm’s time complexity is $O(P\cdot Q)$. Additionally, the poor performance observed when $P=10$ can be attributed to the biased estimation of novelty scores when $P$ is too small, a phenomenon influenced by multiple factors. For a more detailed explanation, please refer to Appendix A.2.
The right panel of Figure 3 demonstrates that when $P$ remains constant, both excessively small and large values of $Q$ negatively impact the algorithm’s performance. This is due to the inaccuracy in local density estimation when $Q$ is too small and the significant reduction in algorithm sensitivity when $Q$ is too large. For further details, please refer to Appendix A.3.
Figure 3: Comparison of AUROC of RND algorithm with different parameters. left: AUROC with different P value when Q=50. right: AUROC with different Q value when P=100
4.3.2 Sensitivity of Design
In our Relative Neighbor Density algorithm, the notion "relative", i.e. comparing idea’s local density with its neighbor’s local densities, plays an important role. Moreover, other distance metric, such as Euclidean distance, could also be used in our algorithm. To understand the sensitivity of the current design, we conducted experiments by changing the design of the "relative" notion, and distance metric. The result presented in Table 3 validate our statement.
Table 3: AUROC Comparison for Different Design . Absolute Local Density: Use the idea’s local density as novelty score. (density calculated by mean distances between idea and idea’s $P$ first level neighbors). Euclidean distance: replace the cosine distance with Euclidean in RND
Test set AUROC
Ours Absolute Local Density Euclidean distance
NeurIPS 0.820 0.851 0.815
Nature Medicine 0.765 0.757 0.753
Mixed 0.795 0.395 0.78
Figure 4: Neighbor Distribution of a Non-novel Idea in Embedding Space (t-SNE processed).
Figure 5: Neighbor Distribution of a Novel Idea in Embedding Space (t-SNE processed).
4.4 Case Study
To visualize the geometric structure of ideas and their neighboring points in the embedding space, we selected two ideas and plotted their surrounding neighbors on a 2-D plane using t-SNE (Van der Maaten and Hinton (2008)).
We first examine Attention is All You Need (Vaswani et al. (2017)), a highly cited article, as an example of a non-novel idea from the current perspective. As shown in Figure 4, there is a dense cluster of neighbors surrounding the idea, while the neighbors around its $P$ nearest neighbors are relatively fewer and more sparsely distributed.
Next, we analyze Evaluating the World Model Implicit in a Generative Model (Vafa et al. (2025)), an article recognized as highly novel by the NeurIPS 2024 Program Chairs based on their comments (openreview (2025)). In Figure 5, the idea’s local neighborhood appears considerably sparser than that of its $P$ nearest neighbors.
Moreover, the differences between the two figures extend beyond the local density around the central idea; the local densities of neighboring ideas also exhibit distinct patterns. In Figure 4, multiple neighboring clusters centered around the idea’s $P$ nearest neighbors are visible, whereas such clustering patterns are absent in the novel idea’s neighborhood. These observations suggest that an idea’s novelty is associated not only with its own local density but also with the density distribution of its neighboring ideas.
5 Discussion
In this work, we proposed a novel neighbor density-based metric for assessing research idea novelty, addressing the limitations of LLM judgment and absolute local density-based metrics. By leveraging large-scale literature embeddings from both biomedical sciences and computer science, our approach ensures robust reliability and cross-domain generalizability. Additionally, we introduced a scalable validation framework that eliminates reliance on expert labeling, enabling objective and reproducible novelty assessment.
5.1 Why a Non-LLM Novelty Assessment Algorithm is Necessary?
While LLMs have the potential to assess novelty, the reliability of their judgment is limited, as outlined in the Introduction section (1). Our experiments (see Table 2) echoed such concern: without an integrated search tool, even the most advanced reasoning models’ performance was comparable to random guessing (AUROC=̃0.5). When a search tool was introduced, Sonnet-3.7 achieved similar accuracy on the NeurIPS test set (AUROC=̃0.8) but experienced significant degradation (AUROC=̃0.6) on both the Nature Medicine and cross-domain test sets.
In contrast, our proposed RND algorithm can produce more reliable and consistent results, as seen in Table 2. Additionally, our algorithm is better at distinguishing genuinely novel ideas from the large pool of candidates from mixing research domains (AUROC=̃0.78 v.s Other’s AUROC<=0.6). Such cross-domain novelty assessment capability is crucial to AI scientist, as more and more innovation happens in the intersection of research domains.
5.2 Domain-invariant Accuracy
The most significant advantage of RND algorithm is in being able to accurately compare novelty across different domains.
As demonstrated in the Table 3, though local density-based HD algorithm and LLM with literature search tools achieved 0.8+ AUROC in NeruIPS test set, their performance degraded in Nature Medicine test set (0.699 for HD, 0.673 for Deepseek-r1). Moreover, when tested in the cross-domain test set (the Mixed test set), the AUROC of HD and LLM significantly degraded below 0.6. In contrast, the AUROC of our proposed RND algorithm remained robust (0.820, 0.765, 0.795) across each single domain and mixed domain test sets.
We argue that such benefits come from RND algorithm’s domain-invariant property, i.e. the distribution of novelty scores produced by RND is identical regardless of the tested domain, which explained why our relative density-based approach succeeds in cross-domain scenarios. According to the mathematical reasoning in Appendix A.4, we concluded the distribution of novelty score $S$ is only subject to $P$ (the number of neighbors considered); thus it is invariant to the validation domain. Furthermore, in Figure 2 (right panel) and 6, the actual distribution of scores echoed the theoretical analysis. Such domain-invariant property is crucial for conducting multi-disciplinary scientific research, where ideas from diverse fields must be compared and evaluated effectively.
5.3 Limitations & Future Work
First, the RND algorithm’s effectiveness is inherently linked to the breadth and quality of the underlying literature database and semantic embedding model. Incomplete literature coverage or suboptimal embedding quality can impact novelty assessment accuracy. While implementing RND requires more infrastructure than LLM-as-judge approaches, this investment enables consistent, deterministic assessments that remain stable across model versions and prompting strategies.
Second, in the test set created by our validation methodology, while the need for expert labeling was avoided, the non-novel samples may be too easily distinguishable from novel ones. By using historical highly-cited papers as non-novel examples, rather than borderline cases such as recently rejected papers or incremental work from current journals, we created a simpler assessment scenario than the actual scenario in real research settings. However, the fact that none of the tested algorithms achieved saturated AUROCs even in this relatively straightforward scenario demonstrates the fundamental challenge of novelty assessment.
While acknowledging these limitations, we believe our work establishes a solid foundation for several promising research directions that could significantly advance automated scientific discovery.
1. Integration with AI Research Workflows: Incorporating our novelty evaluation algorithm into end-to-end AI scientist workflows would enable autonomous research ideation and evaluation. This integration would allow AI systems to independently generate research hypotheses, assess their novelty using our domain-invariant RND algorithm, and prioritize the most promising directions for further investigation. Such integration could accelerate scientific discovery by efficiently navigating complex multi-disciplinary research landscapes where human intuition about novelty is often limited.
2. Enhancing Reasoning Capability for LLM: We propose utilizing our RND algorithm as a sophisticated reward mechanism within reinforcement learning frameworks for training reasoning models for AI research. By providing domain-invariant novelty signals during training, it could potentially guide models to generate more innovative scientific ideas while maintaining scientific validity.
Appendix
Appendix A Algorithms
A.1 Pseudocode
Algorithm 1 Find Neighbors and Calculate Neighbor Density
1:function Neighbor(Input, P, Q)
2:  $\mathbf{v}_{Input}\leftarrow$ Get_Embedding(Input) $\triangleright$ Using M3-Embedding model
3:  C $\leftarrow$ []
4:  neighbors $\leftarrow$ Get_Neighbors($\mathbf{v}_{Input}$, $\max(P,Q)$) $\triangleright$ Find $\max(P,Q)$ nearest neighbors
5:  $neighbors\_for\_count\leftarrow neighbors[:Q]$ $\triangleright$ Only use the Q nearest neighbors to calculate density
6:  $neighbors\_for\_distribution\leftarrow neighbors[:P]$ $\triangleright$ The P nearest neighbors are used to calculate distribution
7:  for each paper in neighbors_for_countneighbors\_for\_count do
8:   $\mathbf{v}_{paper}\leftarrow$ Get_Embedding(paper)
9:   distance $\leftarrow$ 1 - cosine_similarity($\mathbf{v}_{Input},\mathbf{v}_{paper}$)
10:   C.Append(distance)
11:  end for
12:  $ND_{Input}\leftarrow$ mean(C)
13:  return $ND_{Input}$, neighbors_for_distributionneighbors\_for\_distribution
14:end function
Algorithm 2 Calculate Novelty Score of Given Idea
1:Input: Idea
2:Output: A score in the range of 0 to 100
3:D $\leftarrow$ []
[…]
5:for paper in neighbors do
6:  $ND_{paper}$, _ $\leftarrow$ Neighbor(paper, P, Q)
7:  D.Append($ND_{paper}$)
8:end for
9:score $\leftarrow\frac{|\{ND\in D\mid ND\leq ND_{Idea}\}|}{|D|}\times 100$
10:Return score
A.2 Effects of Parameter P
The novelty score is computed as
$\text{score}_{i}=\frac{\left|\{ND\in S_{i}\mid ND\leq ND_{i}\}\right|}{P}\times 100,$ (12)
where $S_{i}=\{ND_{j}\mid j=1,2,\dots,P\}$ is the set of neighbor densities for the $P$ nearest neighbors of the idea $i$, and $ND_{i}$ is the neighbor density for idea $i$ itself.
1. Empirical Cumulative Distribution Function (ECDF) Interpretation
Define the empirical cumulative distribution function (ECDF) for the set $S_{i}$ as
$F_{P}(x)=\frac{1}{P}\sum_{j=1}^{P}\mathbf{1}_{\{ND_{j}\leq x\}},$
where $\mathbf{1}_{\{ND_{j}\leq x\}}$ is the indicator function that is 1 if $ND_{j}\leq x$ and 0 otherwise. Then, by definition, the novelty score can be written as
$\text{score}_{i}=100\cdot F_{P}(ND_{i}).$ (13)
2. Consistency of the ECDF
Let $F(x)$ be the true cumulative distribution function of the neighbor densities (assumed to be i.i.d. samples from a distribution $F$). By the Glivenko-Cantelli theorem, the ECDF $F_{P}(x)$ converges uniformly to $F(x)$ as $P\to\infty$:
$\sup_{x}\left|F_{P}(x)-F(x)\right|\xrightarrow{P\to\infty}0.$
Thus, for a sufficiently large $P$, we have
$F_{P}(ND_{i})\approx F(ND_{i}).$ (14)
This shows that the score, being proportional to $F_{P}(ND_{i})$, converges to $100\cdot F(ND_{i})$, which is the true quantile of $ND_{i}$ in the distribution of neighbor densities.
3. Variance and Sensitivity with Finite $P$
For a finite sample size $P$, $F_{P}(ND_{i})$ is a random variable whose variance depends on $P$. Under the assumption of i.i.d. sampling,
$\operatorname{Var}(F_{P}(ND_{i}))=\frac{F(ND_{i})(1-F(ND_{i}))}{P}.$
Thus, the standard deviation is proportional to $\frac{1}{\sqrt{P}}$. This quantifies that:
  • $\textbf{Smaller }P:$ The variance $\operatorname{Var}(F_{P}(ND_{i}))$ is larger, leading to a noisier (less reliable) estimation of the quantile, and hence of the novelty score.
  • $\textbf{Larger }P:$ The variance decreases, yielding a more accurate estimation of the true quantile $F(ND_{i})$.
The novelty score becomes less sensitive to random fluctuations when $P$ is large, as the empirical quantile is a better estimator of the true quantile.
4. Discreteness of the Score for Small $P$
When $P$ is small, the possible values of $F_{P}(ND_{i})$ are discrete, specifically:
$F_{P}(ND_{i})\in\left\{0,\frac{1}{P},\frac{2}{P},\dots,1\right\}.$
For instance, if $P=1$, then $F_{1}(ND_{i})$ can only be 0 or 1, corresponding to a score of either $0\%$ or $100\%$. This coarse granularity can result in a biased or uninformative measure of novelty. As $P$ increases, the steps $\frac{1}{P}$ become finer, allowing the score to capture more subtle differences in the density distribution.
Conclusion
The parameter $P$ affects the final novelty score in two major ways:
  1. 1.
    Accuracy: As $P$ increases, the empirical cumulative distribution $F_{P}(x)$ better approximates the true cumulative distribution $F(x)$, leading to a more accurate quantile estimate $F(ND_{i})$.
  2. 2.
    Variance: The variance of the estimate $F_{P}(ND_{i})$ is proportional to $\frac{1}{P}$. Thus, a larger $P$ reduces the variability of the score, making it less sensitive to random noise.
In summary, a higher $P$ leads to a more robust and sensitive measure of novelty, while a smaller $P$ results in a discrete and noisier estimate.
A.3 Effects of Parameter Q
The neighbor density (ND) is given by:
$ND=\frac{1}{Q}\sum_{k=1}^{Q}d_{k},$ (15)
where $d_{k}=d(\mathbf{v},\mathbf{v}_{k})$ represents the distance between the point $\mathbf{v}$ and its $k$-th nearest neighbor.
Assuming that $d_{k}$ are independent and identically distributed (i.i.d.) random variables with mean $\mathbb{E}[d_{k}]=\mu_{d}$, we compute the expectation of ND:
$\mathbb{E}[ND]=\mathbb{E}\left[\frac{1}{Q}\sum_{k=1}^{Q}d_{k}\right]=\frac{1}{Q}\sum_{k=1}^{Q}\mathbb{E}[d_{k}]=\frac{1}{Q}Q\mu_{d}=\mu_{d}.$ (16)
The variance of ND is given by:
$\text{Var}(ND)=\text{Var}\left(\frac{1}{Q}\sum_{k=1}^{Q}d_{k}\right).$ (17)
Using the property that the variance of the mean of $Q$ i.i.d. random variables is:
$\text{Var}\left(\frac{1}{Q}\sum_{k=1}^{Q}d_{k}\right)=\frac{1}{Q^{2}}\sum_{k=1}^{Q}\text{Var}(d_{k}).$ (18)
Since each $d_{k}$ has variance $\sigma_{d}^{2}$, we obtain:
$\text{Var}(ND)=\frac{Q\sigma_{d}^{2}}{Q^{2}}=\frac{\sigma_{d}^{2}}{Q}.$ (19)
A.3.1 Interpretation of Variance Scaling
The derived formula:
$\text{Var}(ND)=\frac{\sigma_{d}^{2}}{Q}$ (20)
shows that:
  • As $Q$ increases, the variance of ND decreases.
  • Specifically, variance scales inversely with $Q$, meaning that larger $Q$ results in a more stable estimate of ND.
  • When $Q\to\infty$, $\text{Var}(ND)\to 0$, indicating that ND converges to its expected value $\mu_{d}$, which would cause lost of information on local density.
  • For small $Q$, ND exhibits higher variability, making it more sensitive to local fluctuations.
A.4 Domain-Invariant
A.4.1 Theoretical Analysis
Consider a test set in which each idea is assigned a neighborhood density defined as
$ND=\frac{1}{Q}\sum_{i=1}^{Q}d(\text{idea},a_{i}),$ (21)
where $a_{i}$ denotes the $i$th nearest article in the literature corpus and $d(\cdot,\cdot)$ is the cosine distance.
Let $F(x)$ be the cumulative distribution function (CDF) of the neighborhood densities in the literature corpus. The percentile score for an idea is then defined by
$S=F(ND).$ (22)
By the probability integral transform, if $ND$ is drawn from a distribution with CDF $F(x)$, then
$S\sim\mathcal{U}(0,1).$ (23)
In practice, $F(x)$ is estimated empirically using $P$ articles from the neighborhood. The empirical CDF is given by
$\hat{F}_{P}(x)=\frac{1}{P}\sum_{j=1}^{P}\mathbf{1}\{ND_{j}\leq x\},$ (24)
where $ND_{j}$ is the neighborhood density of the $j$th article, and $\mathbf{1}\{\cdot\}$ is the indicator function. Since
$\sum_{j=1}^{P}\mathbf{1}\{ND_{j}\leq x\}\sim\operatorname{Binomial}(P,F(x)),$ (25)
we have
$\mathbb{E}[\hat{F}_{P}(x)]=F(x)\quad\text{and}\quad\operatorname{Var}[\hat{F}_{P}(x)]=\frac{F(x)(1-F(x))}{P}.$ (26)
Now, consider two literature corpora: a medical corpus with density distribution $F_{M}(x)$ and a computer science corpus with density distribution $F_{C}(x)$. For an idea in the test set, define its scores as
$\hat{S}_{M}=\hat{F}_{M}(ND_{M})\quad\text{and}\quad\hat{S}_{C}=\hat{F}_{C}(ND_{C}),$ (27)
where $ND_{M}$ and $ND_{C}$ are the neighborhood densities computed using the respective corpora. According to equation 26, we have
$\mathbb{E}[\hat{S}]=F(ND)\quad\text{and}\quad\operatorname{Var}[\hat{S}]=\frac{F(ND)(1-F(ND))}{P}.$ (28)
where $F(ND)\sim\mathcal{U}(0,1)$, which implies
$\hat{S}_{M}\stackrel{{\scriptstyle d}}{{=}}\hat{S}_{C}.$ (29)
Furthermore, note that the variance of the empirical estimate $\hat{F}_{P}(x)$ is solely a function of $P$:
$\operatorname{Var}[\hat{F}_{P}(x)]=\frac{F(x)(1-F(x))}{P}.$ (30)
Thus, when $P$ changes (e.g., $P=50$, 100100, or $500$), the change in variance—and hence the fluctuation in the score—is proportional to $\frac{1}{P}$ and is independent of the corpus. In other words,
$\Delta\operatorname{Var}\propto\frac{1}{P},$ (31)
which holds for both the medical and the computer science datasets.
Therefore, we conclude that:
$\boxed{\hat{S}_{M}\stackrel{{\scriptstyle d}}{{=}}\hat{S}_{C}\quad\text{and}\quad\Delta\hat{S}\propto\frac{1}{P}.}$ (32)
This establishes that the scoring distributions for the test set are identical across corporas, and the effect of changing $P$ on the score variation is equivalent for all datasets.
A.4.2 Experimental Evidence
Figure 6: Score Distribution of RND algorithm with different $P$ value.
Appendix B Test Set
B.1 NeurIPS Test Set
Table 4: Titles of Novel (Positive) and Non-novel (Negative) Papers in NeurIPS Test Set
NeurIPS Test Set
Positive Negative
1. Learning to grok: Emergence of in-context learning and skill composition in modular arithmetic tasks 1. Attention is All you Need
2. Nonlocal Attention Operator: Materializing Hidden Knowledge Towards Interpretable Physics Discovery 2. PyTorch: An Imperative Style, High-Performance Deep Learning Library
3. Emergence of Hidden Capabilities: Exploring Learning Dynamics in Concept Space 3. Language Models are Few-Shot Learners
4. Continual learning with the neural tangent ensemble 4. A Unified Approach to Interpreting Model Predictions
5. Neglected Hessian component explains mysteries in sharpness regularization 5. Inductive Representation Learning on Large Graphs
6. Generalization Analysis for Label-Specific Representation Learning 6. Denoising Diffusion Probabilistic Models
7. The Power of Resets in Online Reinforcement Learning 7. GANs Trained by a Two Time-Scale Update Rule Converge to a Local Nash Equilibrium
8. Paths to Equilibrium in Games 8. PointNet++: Deep Hierarchical Feature Learning on Point Sets in a Metric Space
9. Double-Ended Synthesis Planning with Goal-Constrained Bidirectional Search 9. LightGBM: A Highly Efficient Gradient Boosting Decision Tree
10. Time-Reversal Provides Unsupervised Feedback to LLMs 10. Improved Training of Wasserstein GANs
11. Compositional Generalization Across Distributional Shifts with Sparse Tree Operations 11. Improved Techniques for Training GANs
12. Stable Minima Cannot Overfit in Univariate ReLU Networks: Generalization by Large Step Sizes 12. XLNet: Generalized Autoregressive Pretraining for Language Understanding
13. Rule Extrapolation in Language Modeling: A Study of Compositional Generalization on OOD Prompts 13. Prototypical Networks for Few-shot Learning
14. A generalized neural tangent kernel for surrogate gradient learning 14. Convolutional LSTM Network: A Machine Learning Approach for Precipitation Nowcasting
15. GREATS: Online Selection of High-Quality Data for LLM Training in Every Iteration 15. Convolutional Neural Networks on Graphs with Fast Localized Spectral Filtering
16. Non-Asymptotic Uncertainty Quantification in High-Dimensional Learning 16. Spatial Transformer Networks
17. Boosting Vision-Language Models with Transduction 17. Matching Networks for One Shot Learning
18. Input-to-State Stable Coupled Oscillator Networks for Closed-form Model-based Control in Latent Space 18. Learning both Weights and Connections for Efficient Neural Network
19. Assouad, Fano, and Le Cam with Interaction: A Unifying Lower Bound Framework and Characterization for Bandit Learnability 19. Bootstrap Your Own Latent: A New Approach to Self-Supervised Learning
20. Exploring Jacobian Inexactness in Second-Order Methods for Variational Inequalities: Lower Bounds, Optimal Algorithms and Quasi-Newton Approximations 20. Character-level Convolutional Networks for Text Classification
21. Who’s asking? User personas and the mechanics of latent misalignment 21. R-FCN: Object Detection via Region-based Fully Convolutional Networks
22. Self-Consuming Generative Models with Curated Data Provably Optimize Human Preferences 22. Simple and Scalable Predictive Uncertainty Estimation using Deep Ensembles
23. Selective Generation for Controllable Language Models 23. wav2vec 2.0: A Framework for Self-Supervised Learning of Speech Representations
24. Constrained Adaptive Attack: Effective Adversarial Attack Against Deep Neural Networks for Tabular Data 24. Neural Ordinary Differential Equations
25. Learning Generalized Linear Programming Value Functions 25. Dynamic Routing Between Capsules
26. Optimizing Automatic Differentiation with Deep Reinforcement Learning 26. What Uncertainties Do We Need in Bayesian Deep Learning for Computer Vision?
27. Overcoming Common Flaws in the Evaluation of Selective Classification Systems 27. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks
28. Revisiting K-mer Profile for Effective and Scalable Genome Representation Learning 28. Neural Discrete Representation Learning
29. Trading Place for Space: Increasing Location Resolution Reduces Contextual Capacity in Hippocampal Codes 29. InfoGAN: Interpretable Representation Learning by Information Maximizing Generative Adversarial Nets
30. Reproducibility of predictive networks for mouse visual cortex 30. Multi-Agent Actor-Critic for Mixed Cooperative-Competitive Environments
31. Nonlinear dynamics of localization in neural receptive fields 31. Supervised Contrastive Learning
32. Learning Noisy Halfspaces with a Margin: Massart is No Harder than Random 32. Equality of Opportunity in Supervised Learning
33. Cracking the Code of Juxtaposition: Can AI Models Understand the Humorous Contradictions 33. Unsupervised Learning of Visual Features by Contrasting Cluster Assignments
34. Evaluating the World Model Implicit in a Generative Model 34. Mean teachers are better role models: Weight-averaged consistency targets improve semi-supervised deep learning results
35. TrackIME: Enhanced Video Point Tracking via Instance Motion Estimation 35. Teaching Machines to Read and Comprehend
36. DiffLight: A Partial Rewards Conditioned Diffusion Model for Traffic Signal Control with Missing Data 36. ViLBERT: Pretraining Task-Agnostic Visiolinguistic Representations for Vision-and-Language Tasks
37. Mean-Field Langevin Dynamics for Signed Measures via a Bilevel Approach 37. Convolutional Networks on Graphs for Learning Molecular Fingerprints
38. Stabilized Proximal-Point Methods for Federated Optimization 38. Generative Modeling by Estimating Gradients of the Data Distribution
39. Reparameterization invariance in approximate Bayesian inference 39. FixMatch: Simplifying Semi-Supervised Learning with Consistency and Confidence
40. Disentangling the Roles of Distinct Cell Classes with Cell-Type Dynamical Systems 40. Learning Structured Output Representation using Deep Conditional Generative Models
41. Linear Regression using Heterogeneous Data Batches 41. Glow: Generative Flow with Invertible 1x1 Convolutions
42. A Near-optimal Algorithm for Learning Margin Halfspaces with Massart Noise 42. CatBoost: unbiased boosting with categorical features
43. Neural Krylov Iteration for Accelerating Linear System Solving 43. Man is to Computer Programmer as Woman is to Homemaker? Debiasing Word Embeddings
44. Human Expertise in Algorithmic Prediction 44. Neural Tangent Kernel: Convergence and Generalization in Neural Networks
45. Analysing Multi-Task Regression via Random Matrix Theory with Application to Time Series Forecasting 45. Generative Adversarial Imitation Learning
46. No-regret Learning in Harmonic Games: Extrapolation in the Face of Conflicting Interests 46. Pointer Networks
47. Learning diffusion at lightspeed 47. BinaryConnect: Training Deep Neural Networks with binary weights during propagations
48. Voila-A: Aligning Vision-Language Models with User’s Gaze Attention 48. MixMatch: A Holistic Approach to Semi-Supervised Learning
49. Barely Random Algorithms and Collective Metrical Task Systems 49. Unsupervised Image-to-Image Translation Networks
50. Goal Reduction with Loop-Removal Accelerates RL and Models Human Brain Activity in Goal-Directed Learning 50. Deep Reinforcement Learning from Human Preferences
51. BricksRL: A Platform for Democratizing Robotics and Reinforcement Learning Research and Education with LEGO 51. Cross-lingual Language Model Pretraining
52. Breaking Long-Tailed Learning Bottlenecks: A Controllable Paradigm with Hypernetwork-Generated Diverse Experts 52. Attention-Based Models for Speech Recognition
53. Kermut: Composite kernel regression for protein variant effects 53. End-To-End Memory Networks
54. Automatically Learning Hybrid Digital Twins of Dynamical Systems 54. Gradient Episodic Memory for Continual Learning
55. On the Identifiability of Poisson Branching Structural Causal Model Using Probability Generating Function 55. Open Graph Benchmark: Datasets for Machine Learning on Graphs
56. Weisfeiler and Leman Go Loopy: A New Hierarchy for Graph Representational Learning 56. Conditional Image Generation with PixelCNN Decoders
57. Unlocking the Capabilities of Thought: A Reasoning Boundary Framework to Quantify and Optimize Chain-of-Thought 57. Generalized Cross Entropy Loss for Training Deep Neural Networks with Noisy Labels
58. Reinforcement Learning Under Latent Dynamics: Toward Statistical and Algorithmic Modularity 58. Skip-Thought Vectors
59. Generalization Error Bounds for Two-stage Recommender Systems with Tree Structure 59. Self-Normalizing Neural Networks
60. Can Transformers Smell Like Humans? 60. PointCNN: Convolution On X-Transformed Points
61. Geodesic Optimization for Predictive Shift Adaptation on EEG data 61. Learning Structured Sparsity in Deep Neural Networks
62. Second-order forward-mode optimization of recurrent neural networks for neuroscience 62. Implicit Neural Representations with Periodic Activation Functions
63. Discrete Flow Matching 63. Deep Generative Image Models using a Laplacian Pyramid of Adversarial Networks
64. Motion Forecasting in Continuous Driving 64. Hindsight Experience Replay
65. Moving Off-the-Grid: Scene-Grounded Video Representations 65. Unsupervised Data Augmentation for Consistency Training
66. Aligner: Efficient Alignment by Learning to Correct 66. Conditional Adversarial Domain Adaptation
67. Questioning the Survey Responses of Large Language Models 67. SuperGLUE: A Stickier Benchmark for General-Purpose Language Understanding Systems
68. Saliency-driven Experience Replay for Continual Learning 68. Big Self-Supervised Models are Strong Semi-Supervised Learners
69. Adversarial Environment Design via Regret-Guided Diffusion Models 69. Fourier Features Let Networks Learn High Frequency Functions in Low Dimensional Domains
70. Localized Zeroth-Order Prompt Optimization 70. Improved Deep Metric Learning with Multi-class N-pair Loss Objective
71. Molecule Design by Latent Prompt Transformer 71. Hierarchical Graph Representation Learning with Differentiable Pooling
72. Can Learned Optimization Make Reinforcement Learning Less Difficult? 72. Scheduled Sampling for Sequence Prediction with Recurrent Neural Networks
73. Reverse Transition Kernel: A Flexible Framework to Accelerate Diffusion Inference 73. Learning to learn by gradient descent by gradient descent
74. Any2Graph: Deep End-To-End Supervised Graph Prediction With An Optimal Transport Loss 74. Co-teaching: Robust training of deep neural networks with extremely noisy labels
75. MInference 1.0: Accelerating Pre-filling for Long-Context LLMs via Dynamic Sparse Attention 75. Continual Learning with Deep Generative Replay
76. Ensemble Learning for Heterogeneous Large Language Models with Deep Parallel Collaboration 76. Learning a Probabilistic Latent Space of Object Shapes via 3D Generative-Adversarial Modeling
77. Humanoid Locomotion as Next Token Prediction 77. Big Bird: Transformers for Longer Sequences
78. NeoRL: Efficient Exploration for Nonepisodic RL 78. Weight Normalization: A Simple Reparameterization to Accelerate Training of Deep Neural Networks
79. Toxicity Detection for Free 79. A Simple Unified Framework for Detecting Out-of-Distribution Samples and Adversarial Attacks
80. Semi-supervised Multi-label Learning with Balanced Binary Angular Margin Loss 80. Sanity Checks for Saliency Maps
81. When Does Label Smoothing Help?
82. Graph Contrastive Learning with Augmentations
83. Link Prediction Based on Graph Neural Networks
84. Improved Variational Inference with Inverse Autoregressive Flow
85. Training Generative Adversarial Networks with Limited Data
86. Visualizing the Loss Landscape of Neural Nets
87. Adversarial Examples Are Not Bugs, They Are Features
88. HiFi-GAN: Generative Adversarial Networks for Efficient and High Fidelity Speech Synthesis
89. Understanding the Effective Receptive Field in Deep Convolutional Neural Networks
90. Learning to summarize from human feedback
91. Federated Multi-Task Learning
92. Training Very Deep Networks
93. Efficient and Robust Automated Machine Learning
94. A Theoretically Grounded Application of Dropout in Recurrent Neural Networks
95. Generating Diverse High-Fidelity Images with VQ-VAE-2
96. f-GAN: Training Generative Neural Samplers using Variational Divergence Minimization
97. Coupled Generative Adversarial Networks
98. Conservative Q-Learning for Offline Reinforcement Learning
99. A simple neural network module for relational reasoning
B.2 Nature Medicine Test Set
Table 5: Titles of Novel (Positive) and Non-novel (Negative) Papers in Nature Medicine Test Set
Nature Medicine Test Set
Positive Negative
1. Sleep patterns and risk of chronic disease as measured by long-term monitoring with commercial wearable devices in the All of Us Research Program. 1. Microenvironmental regulation of tumor progression and metastasis
2. Botensilimab plus balstilimab in relapsed/refractory microsatellite stable metastatic colorectal cancer: a phase 1 trial. 2. Temporal dynamics in viral shedding and transmissibility of COVID-19
3. Lipidome changes due to improved dietary fat quality inform cardiometabolic risk reduction and precision nutrition. 3. The Consensus Molecular Subtypes of Colorectal Cancer
4. Fratricide-resistant CD7-CAR T cells in T-ALL. 4. High-performance medicine: the convergence of human and artificial intelligence
5. International multicenter validation of AI-driven ultrasound detection of ovarian cancer. 5. Understanding the tumor immune microenvironment (TIME) for effective therapy
6. Donor-derived GD2-specific CAR T cells in relapsed or refractory neuroblastoma. 6. Intestinal microbiota metabolism of L-carnitine, a nutrient in red meat, promotes atherosclerosis
7. Single-nucleus chromatin accessibility and transcriptomic map of breast tissues of women of diverse genetic ancestry. 7. Post-acute COVID-19 syndrome
8. Unidirectional association of clonal hematopoiesis with atherosclerosis development. 8. Melanoma exosomes educate bone marrow progenitor cells toward a pro-metastatic phenotype through MET
9. Echocardiographic screening for heart failure and optimization of the care pathway for individuals with pacemakers: a randomized controlled trial. 9. Signatures of T cell dysfunction and exclusion predict cancer immunotherapy response
10. Population-based, first-tier genomic newborn screening in the maternity ward. 10. Neutralizing antibody levels are highly predictive of immune protection from symptomatic SARS-CoV-2 infection
11. Allogeneic CD5-specific CAR-T therapy for relapsed/refractory T-ALL: a phase 1 trial. 11. Ischemia and reperfusion—from mechanism to translation
12. Transplantation of a genetically modified porcine heart into a live human. 12. Mechanisms of fibrosis: therapeutic translation for fibrotic disease
13. A multi-modal single-cell and spatial expression map of metastatic breast cancer biopsies across clinicopathological features. 13. Metabolite profiles and the risk of developing diabetes
14. ctDNA-based molecular residual disease and survival in resectable colorectal cancer. 14. Mechanisms of NAFLD development and therapeutic strategies
15. Antifungal heteroresistance causes prophylaxis failure and facilitates breakthrough Candida parapsilosis infections. 15. Inflammasomes: mechanism of action, role in disease, and therapeutics
16. Subcutaneous weekly semaglutide with automated insulin delivery in type 1 diabetes: a double-blind, randomized, crossover trial. 16. Chronic inflammation in the etiology of disease across the life span
17. Combined endurance and resistance exercise training in heart failure with preserved ejection fraction: a randomized controlled trial. 17. Mutational Landscape of Metastatic Cancer Revealed from Prospective Clinical Sequencing of 10,000 Patients
18. Multi-omic profiling a defined bacterial consortium for treatment of recurrent Clostridioides difficile infection. 18. Antibody responses to SARS-CoV-2 in patients with COVID-19
19. An organotypic atlas of human vascular cells. 19. ABT-199, a potent and selective BCL-2 inhibitor, achieves antitumor activity while sparing platelets
20. Lipid profiling identifies modifiable signatures of cardiometabolic risk in children and adolescents with obesity. 20. Clinical and immunological assessment of asymptomatic SARS-CoV-2 infections
21. Ferric carboxymaltose for anemia in late pregnancy: a randomized controlled trial. 21. Extrapulmonary manifestations of COVID-19
22. Effects of conditional cash transfers on tuberculosis incidence and mortality according to race, ethnicity and socioeconomic factors in the 100 Million Brazilian Cohort. 22. A guide to deep learning in healthcare
23. Phenome-wide associations of sleep characteristics in the Human Phenotype Project. 23. A global survey of potential acceptance of a COVID-19 vaccine
24. Proteomic signatures improve risk prediction for common and rare diseases. 24. The emerging role of lncRNAs in cancer
25. Remotely delivered weight management for people with long COVID and overweight: the randomized wait-list-controlled ReDIRECT trial. 25. SARS-CoV-2 Entry Genes Are Most Highly Expressed in Nasal Goblet and Ciliated Cells within Human Airways
26. Sustained effect of prasinezumab on Parkinson’s disease motor progression in the open-label extension of the PASADENA trial. 26. Gut microbiota metabolism of dietary fiber influences allergic airway disease and hematopoiesis
27. Collaboration between clinicians and vision-language models in radiology report generation. 27. The immunology of stroke: from mechanisms to translation
28. Oral obeldesivir provides postexposure protection against Marburg virus in nonhuman primates. 28. Asthma phenotypes: the evolution from clinical to molecular approaches
29. Digital consults in heart failure care: a randomized controlled trial. 29. Single-cell landscape of bronchoalveolar immune cells in patients with COVID-19
30. Increased frequency of repeat expansion mutations across different populations. 30. A small-molecule inhibitor of the NLRP3 inflammasome for the treatment of inflammatory diseases
31. Autogene cevumeran with or without atezolizumab in advanced solid tumors: a phase 1 trial. 31. Respiratory virus shedding in exhaled breath and efficacy of face masks
32. A high-performance brain-computer interface for finger decoding and quadcopter game control in an individual with paralysis. 32. Cancer stem cells revisited
33. Mapping the effectiveness and risks of GLP-1 receptor agonists. 33. Atherosclerosis: current pathogenesis and therapeutic options
34. Evaluating generalizability of oncology trial results to real-world patients using machine learning-based trial emulations. 34. Brown and beige fat: development, function and therapeutic potential
35. AI-based differential diagnosis of dementia etiologies on multimodal data. 35. Cardiologist-level arrhythmia detection and classification in ambulatory electrocardiograms using a deep neural network
36. Molecular classification to refine surgical and radiotherapeutic decision-making in meningioma. 36. An inflammatory cytokine signature predicts COVID-19 severity and survival
37. A framework for sharing of clinical and genetic data for precision medicine applications. 37. Classification and mutation prediction from non–small cell lung cancer histopathology images using deep learning
38. A generalist medical language model for disease diagnosis assistance. 38. CSF-1R inhibition alters macrophage polarization and blocks glioma progression
39. Subclassification of obesity for precision prediction of cardiometabolic diseases. 39. Clinically applicable deep learning for diagnosis and referral in retinal disease
40. Somatic CAG repeat expansion in blood associates with biomarkers of neurodegeneration in Huntington’s disease decades before clinical motor diagnosis. 40. Development, maintenance and disruption of the blood-brain barrier
41. Genomic reanalysis of a pan-European rare-disease resource yields new diagnoses. 41. Adipocytes promote ovarian cancer metastasis and provide energy for rapid tumor growth
42. In vivo base editing extends lifespan of a humanized mouse model of prion disease. 42. An ultrasensitive method for quantitating circulating tumor DNA with broad patient coverage
43. Self-improving generative foundation model for synthetic medical image generation and clinical applications. 43. The cancer stem cell: premises, promises and challenges
44. Data-driven cluster analysis identifies distinct types of metabolic dysfunction-associated steatotic liver disease. 44. Attributes and predictors of long COVID
45. The economic value of reducing avoidable mortality. 45. WNT signaling in bone homeostasis and disease: from human mutations to treatments
46. Genetic basis of early onset and progression of type 2 diabetes in South Asians. 46. The role of autophagy in neurodegenerative disease
47. Posthospitalization COVID-19 cognitive deficits at 1 year are global and associated with elevated brain injury markers and gray matter volume reduction. 47. Clinical-grade computational pathology using weakly supervised deep learning on whole slide images
48. Safety and reactogenicity of a controlled human infection model of sand fly-transmitted cutaneous leishmaniasis. 48. A human memory T-cell subset with stem cell-like properties
49. Cabozantinib and nivolumab with or without live bacterial supplementation in metastatic renal cell carcinoma: a randomized phase 1 trial. 49. Current understanding of the human microbiome
50. Seven-year performance of a clinical metagenomic next-generation sequencing test for diagnosis of central nervous system infections. 50. Molecular analysis of gastric cancer identifies subtypes associated with distinct clinical outcomes
51. Brain aging patterns in a large and diverse cohort of 49,482 individuals. 51. Cellular senescence in aging and age-related disease: from mechanisms to therapy
[…]
53. Cytokine-mediated CAR T therapy resistance in AML. 53. Resting-state connectivity biomarkers define neurophysiological subtypes of depression
54. Prediction of brain metastasis development with DNA methylation signatures. 54. Age-dependent effects in the transmission and control of COVID-19 epidemics
55. Personalized, autologous neoantigen-specific T cell therapy in metastatic melanoma: a phase 1 trial. 55. Tumor angiogenesis: molecular pathways and therapeutic targets
56. A generalist vision-language foundation model for diverse biomedical tasks. 56. Evidence for osteocyte regulation of bone homeostasis through RANKL expression
57. DNA liquid biopsy-based prediction of cancer-associated venous thromboembolism. 57. Modelling the COVID-19 epidemic and implementation of population-wide interventions in Italy
58. Semaglutide in patients with overweight or obesity and chronic kidney disease without diabetes: a randomized double-blind placebo-controlled clinical trial. 58. Identification of the molecular basis of doxorubicin-induced cardiotoxicity
59. Intracerebroventricular B7-H3-targeting CAR T cells for diffuse intrinsic pontine glioma: a phase 1 trial. 59. New from NPG: Genome-wide association study identifies five new schizophrenia loci
60. AI-based selection of individuals for supplemental MRI in population-based breast cancer screening: the randomized ScreenTrustMRI trial. 60. Senolytics Improve Physical Function and Increase Lifespan in Old Age
61. A toolbox for surfacing health equity harms and biases in large language models. 61. Subtypes of Pancreatic Ductal Adenocarcinoma and Their Differing Responses to Therapy
62. Partitioned polygenic risk scores identify distinct types of metabolic dysfunction-associated steatotic liver disease. 62. A purified membrane protein from Akkermansia muciniphila or the pasteurized bacterium improves metabolism in obese and diabetic mice
63. Multi-omics-based mapping of decidualization resistance in patients with a history of severe preeclampsia. 63. The NALP3/NLRP3 Inflammasome Instigates Obesity-Induced Autoinflammation and Insulin Resistance
64. Electronic nudges for sustained influenza vaccination uptake in older adults: the nationwide randomized NUDGE-FLU-2 trial. 64. IgE and mast cells in allergic disease
65. A time-stratified, case-crossover study of heat exposure and perinatal mortality from 16 hospitals in sub-Saharan Africa. 65. Brown adipose tissue activity controls triglyceride clearance
66. SARS-CoV-2 correlates of protection from infection against variants of concern. 66. Intraoperative tumor-specific fluorescence imaging in ovarian cancer by folate receptor-$\alpha$ targeting: first in-human results
67. The cellular and signaling networks linking the immune system and metabolism in disease
68. Supplementation with Akkermansia muciniphila in overweight and obese human volunteers: a proof-of-concept exploratory study
69. Why don’t we get more cancer? A proposed role of the microenvironment in restraining cancer progression
70. Clearance of senescent cells by ABT263 rejuvenates aged hematopoietic stem cells in mice
71. 4-1BB Costimulation Ameliorates T Cell Exhaustion Induced by Tonic Signaling of Chimeric Antigen Receptors
72. Characteristics of pediatric SARS-CoV-2 infection and potential evidence for persistent fecal viral shedding
73. Molecular subtypes of diffuse large B cell lymphoma are associated with distinct pathogenic mechanisms and outcomes
74. The oral and gut microbiomes are perturbed in rheumatoid arthritis and partly normalized after treatment
75. End-to-end lung cancer screening with three-dimensional deep learning on low-dose chest computed tomography
76. The practical implementation of artificial intelligence technologies in medicine
77. Estimating clinical severity of COVID-19 from the transmission dynamics in Wuhan, China
78. Microglia emerge as central players in brain disease
79. Organ reengineering through development of a transplantable recellularized liver graft using decellularized liver matrix
80. RET, ROS1 and ALK fusions in lung cancer
81. Divergent clonal evolution of castration resistant neuroendocrine prostate cancer
82. Long-term cardiovascular outcomes of COVID-19
83. Ketone body $\beta$-hydroxybutyrate blocks the NLRP3 inflammasome-mediated inflammatory disease
84. Large language models in medicine
85. In vivo photodynamic therapy using upconversion nanoparticles as remote-controlled nanotransducers
86. Mitochondrial transfer from bone-marrow–derived stromal cells to pulmonary alveoli protects against acute lung injury
87. A single-cell atlas of the peripheral immune response in patients with severe COVID-19
88. Determinants of response and resistance to CD19 chimeric antigen receptor (CAR) T cell therapy of chronic lymphocytic leukemia
89. Cancer epigenetics reaches mainstream oncology
90. Real-time tracking of self-reported symptoms to predict potential COVID-19
91. Metformin alters the gut microbiome of individuals with treatment-naive type 2 diabetes, contributing to the therapeutic effects of the drug
92. Synaptic plasticity and depression: new insights from stress and rapid-acting antidepressants
93. Matrix-embedded cells control osteoclast formation
94. Targeting EZH2 in cancer
95. Comprehensive molecular characterization of clinical responses to PD-1 inhibition in metastatic gastric cancer
96. Identification of miR-34a as a potent inhibitor of prostate cancer progenitor cells and metastasis by directly repressing CD44
97. Phenotype molding of stromal cells in the lung tumor microenvironment
98. Key roles of adjuvants in modern vaccines
99. AI in health and medicine
Appendix C Prompt
C.1 Prompt for LLM with NeurIPS 2024 Review Guideline
Table 6: Prompt for LLM with NeurIPS 2024 Review Guideline
Prompt
Task description: You are a researcher who is reviewing a paper that was submitted to a computer science venue. Be critical and cautious in your decision. If a paper is bad or you are unsure, give it bad scores and reject it. Below is a description of the questions you will be asked on the review form for each paper and some guidelines on what to consider when answering these questions.
Reviewer guidelines: 1. Summary: Briefly summarize the paper and its contributions. This is not the place to critique the paper; the authors should generally agree with a well-written summary.
2. Strengths and Weaknesses: Please provide a thorough assessment of the strengths and weaknesses of the paper, touching on each of the following dimensions:
- Originality: Are the tasks or methods new? Is the work a novel combination of well-known techniques? (This can be valuable!) Is it clear how this work differs from previous contributions?
- Quality: Is the submission technically sound? Are claims well-supported (e.g., by theoretical analysis or experimental results)? Are the methods used appropriately? Is this a complete piece of work or a work in progress? Are the authors careful and honest about evaluating both the strengths and weaknesses of their work?
- Clarity: Is the submission clearly written? Is it well organized? (If not, please make constructive suggestions for improving its clarity.) Does it adequately inform the reader? (Note that a superbly written paper provides enough information for an expert reader to reproduce its results.)
- Significance: Are the results important? Are others (researchers or practitioners) likely to use the ideas or build on them? Does the submission address a difficult task in a better way than previous work? Does it advance the state of the art in a demonstrable way? Does it provide unique data, unique conclusions about existing data, or a unique theoretical or experimental approach?
3. Questions: Please list and carefully describe any questions and suggestions for the authors. Think of the things where a response from the author can change your opinion, clarify confusion, or address a limitation. This can be very important for a productive rebuttal and discussion phase with the authors.
4. Ethical concerns: If there are ethical issues with this paper, please flag the paper for an ethics review.
5. Overall: Please provide an "overall score" for this submission.
Choices:
- 10: Award quality: Technically flawless paper with groundbreaking impact on one or more areas, with exceptionally strong evaluation, reproducibility, and resources, and no unaddressed ethical considerations.
- 9: Very Strong Accept: Technically flawless paper with groundbreaking impact on at least one area and excellent impact on multiple areas, with flawless evaluation, resources, and reproducibility, and no unaddressed ethical considerations.
- 8: Strong Accept: Technically strong paper, with novel ideas, excellent impact on at least one area or high-toexcellent impact on multiple areas, with excellent evaluation, resources, and reproducibility, and no unaddressed ethical considerations.
- 7: Accept: Technically solid paper, with high impact on at least one sub-area or moderate-to-high impact on more than one area, with good-to-excellent evaluation, resources, reproducibility, and no unaddressed ethical considerations.
- 6: Weak Accept: Technically solid, moderate-to-high impact paper, with no major concerns with respect to evaluation, resources, reproducibility, and ethical considerations.
- 5: Borderline accept: Technically solid paper where reasons to accept outweigh reasons to reject, e.g., limited evaluation. Please use sparingly.
- 4: Borderline reject: Technically solid paper where reasons to reject, e.g., limited evaluation, outweigh reasons to accept, e.g., good evaluation. Please use sparingly.
- 3: Reject: For instance, a paper with technical flaws, weak evaluation, inadequate reproducibility, and incompletely addressed ethical considerations.
- 2: Strong Reject: For instance, a paper with major technical flaws, and/or poor evaluation, limited impact, poor reproducibility, and mostly unaddressed ethical considerations.
- 1: Very Strong Reject: For instance, a paper with trivial results or unaddressed ethical considerations
Provided paper: Here is the paper you are asked to review:
{paper} Output:
Return a JSON object:
<JSON>
template
<JSON>
C.2 Prompt for Standardized Project Proposals
Table 7: Prompt for Standardized Project Proposals
Prompt
Role: You are a writing assistant specialized in editing academic writing.
Task: I will give you a student’s research idea and an idea template. Your task is to edit the student’s idea to follow the template’s format.
Student idea:
Title
{title}
Main Idea
{paper}
Template:
1. Title: A concise statement of the main research question to be used as the paper title.
2. Problem Statement: Clearly define the problem your research intends to address. Explain clearly why this problem is interesting and important.
3. Motivation: Explain why existing methods are not good enough to solve the problem, and explain the inspiration behind the new proposed method. You should also motivate why the proposed method would work better than existing baselines on the problem.
4. Proposed Method: Explain how the proposed method works, describe all the essential steps.
5. Step-by-Step Experiment Plan: Break down every single step of the experiments, make sure every step is executable. Cover all essential details such as the datasets, models, and metrics to be used. If the project involves prompting, give some example prompts for each step.
6. Test Case Examples: Give at least two concrete examples. The first example should show how the baseline method fails on the test case. If there are multiple baselines, give examples for all of them. The second example should show how the proposed method succeeds on the test case. For each test case, include the input (test example and the full prompt) and the expected output. You should also provide an explanation for why the outputs from the proposed prompt are better. If the proposed method has multiple steps, break them down into intermediate steps.
7. Fallback Plan: Propose some alternative plans for what should the students do if the proposed method doesn’t manage to satisfy the success criteria. For example, you can suggest additional analysis to help debug why the proposed method didn’t work, which could inform alternative new methods, or just turn the project into an analysis paper instead by offering some interesting ablation and insights.
Requirement:
Make sure that you only edit the wording and formatting, including things like punctuation, capitalization, linebreaks, and bullet points. Also make sure to edit any informal wording and phrasing to use vocabulary that sounds like the template’s writing style. No other changes are allowed beyond these.
You should use tab as indentation and make sure to use appropriate nested indentation for sub-bullets. All bullets should have a clear hierarchy so people can easily differentiate the sub-bullets. Only leave empty lines between sections and remove any extra line breaks. If many bullet points are clustered together in a paragraph, separate them clearly with indentation and appropriate bullet point markers. Change to a new line for each new bullet point.
For the fallback plan, do not list a bunch of bullet points. Instead, condense them into one coherent paragraph. For line breaks, avoid Raw String Literals or Double Backslashes when using "
n", and change them to spaces or tabs.
For in-line citations, if the citation mentioned the author’s last name (like "(Si et al., 2023)" or "(An et al., 2024)"), you should keep them there; but if the citation is just a number (like "[1]" or "[3,4,5]"), you should just remove it and do some necessary rephrasing to make the sentence still sound coherent without the references.
Apart from minor rephrasing and changing formatting, do not change any content of the idea. You must preserve the exact meaning of the original idea, do not change, remove, or add any other details. Do not drop any sections (including test case examples). Do not rename any models, datasets, or methods. Do not drop clarification or examples in brackets and do not drop any data source mentions (e.g., Chatbot Arena or Wildchat)! Note that when indexing test case examples, each test case example could have multiple steps of inputs and outputs and you shouldn’t give separate indices to them. Each test case example should be a whole set of input-output pairs for the baseline(s) and proposed method.
For the proposed method section, avoid any big changes. If the section comes in as a coherent paragraph, you don’t have to break it down into bullet points. If the section is already in bullet points, you should keep it that way. If the section is a mix of both, you should keep the bullet points and the coherent paragraph as they are. Keep all the clarification and examples mentioned in all the sections and do not remove any of them (including those in brackets).
For model selection, if any version of Claude is mentioned, change it to the latest version of Claude (Claude-3.5); if any version of LLaMA is mentioned, change it to the latest version LLaMA-3. Do not make any other model changes. Now directly generate the edited student idea to match the format of the template.
C.3 Prompt for LLM with Literature Search
Table 8: Prompt for LLM with Literature Search
Prompt
Role: You are an ambitious AI PhD student who is looking to publish a paper that will contribute significantly to the field.
Task description:
You have an idea and you want to check if it is novel or not. I.e., not overlapping significantly with existing literature or already well explored. Be a harsh critic for novelty, ensure there is a sufficient contribution in the idea for a new conference or workshop paper.
You will be given the titles and abstracts of the 10 papers most relevant to your idea. Decide a paper idea is novel if after sufficient searching, you have not found a paper that significantly overlaps with your idea. Decide a paper idea is not novel, if you have found a paper that significantly overlaps with your idea.
Set your decision to True if you think the idea is novel, set it to False if you think the idea is not novel.
Your Idea:
This is the idea you need to judge for novelty:
{Idea}
Top 10 relevant papers:
{papers}
Output:
Return only True or False, dont return any other words.
Appendix D Result in Detail
D.1 AUROC of Different LLM-Based Model in Detail
Sonnet-3.7 with guideline Sonnet-3.7 with tournament LLM + literature search
Sonnet-3.7 Deepseek-r1 Gpt-4o
Round1 NeurIPS 0.544 0.497 0.818 0.746 0.56
Nature Medicine NaN 0.501 0.616 0.663 0.518
Mixed NaN 0.5 0.596 0.583 0.492
Round2 NeurIPS 0.505 0.496 0.799 0.701 0.578
Nature Medicine NaN 0.51 0.624 0.707 0.551
Mixed NaN 0.497 0.603 0.661 0.535
Round3 NeurIPS 0.59 0.496 0.823 0.682 0.564
Nature Medicine NaN 0.497 0.609 0.649 0.566
Mixed NaN 0.506 0.593 0.543 0.54
Table 9: AUROC of Different LLM-Based Model in Detail
D.2 Comparison of AUROC of RND algorithm with different P values
Test set AUROC
P=10,Q=50 P=50,Q=50 P=100,Q=50 P=500,Q=50 P=1000,Q=50
NeurIPS 0.808 0.830 0.820 0.825 0.833
Nature Medicine 0.686 0.754 0.765 0.760 0.765
Mixed 0.762 0.805 0.795 0.786 0.787
Table 10: Comparison of AUROC of RND algorithm with different P values
D.3 Comparison of AUROC of RND algorithm with different Q values
Test set AUROC
P=100,Q=2 P=100,Q=5 P=100,Q=10 P=100,Q=50 P=100,Q=100 P=100,Q=500 P=100,Q=1000
NeurIPS 0.778 0.818 0.827 0.820 0.808 0.788 0.768
Nature Medicine 0.603 0.686 0.723 0.765 0.757 0.744 0.722
Mixed 0.628 0.745 0.774 0.795 0.782 0.771 0.758
Table 11: Comparison of AUROC of RND algorithm with different Q values
Appendix E Computing Resources
To conduct our experiment, the necessary computing resources includes
1. Literature database with embedding vector: we used a server with 16 CPU cores, 128GB RAM, 2TB SSD disk(which actually used around 1TB) to deploy an ElasticSearch version 8 as literature search engine. The embedding of each literature is obtained using public available API services.
2. Algorithm development and experiment: a server with 16 CPU cores, 128GB RAM.